find the roots using bisector. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

e^x+x=4
1
Expert's answer
2020-10-26T20:10:49-0400
"f(x)=e^x+x-4"

"a=0, b=2"

"f(a)=f(0)=e^0 +0-4=-3"

"f(b)=f(2)=e^2 +2-4=e^2-2"

"f(a)f(b)=f(0)f(2)=-3(e^2-2)<0"


"x_n=\dfrac{a_n+b_n}{2}"

"a_{n+1}=x_n, b_{n+1}=b_n, f(a_n)f(x_n)\geq0"

"a_{n+1}=a_n, b_{n+1}=x_n, f(b_n)f(x_n)\geq0"

"|f(x_n)|\leq \varepsilon=>answer =x_n"


"\begin{matrix}
 n & x_n & f(x_n)\\
 0 & 1 & -0.28171817 \\
 1 & 1.5 & 1.9868907 \\
 2 & 1.25 & 0.74034296 \\
 3 & 1.125 & 0.20521685 \\
 4 & 1.0625 & -0.04390406 \\
 5 & 1.09375 & 0.07919854 \\
 6 & 1.078125 & 0.01728845 \\
 7 & 1.0703125 & -0.01339680 \\
 8 & 1.07421875 & 0.00192349 \\
 9 & 1.072265625 & -0.00574223 \\
 10 & 1.0732421875 & -0.00191077 \\
 11 & 1.07373046875 & 0.00000601 \\
 
\end{matrix}"


"\big|\dfrac{1.07373046875-1.0732421875}{1.0732421875}\big|\cdot100\%\approx"

"\approx0.0455\%<0.05\%"

Root is "1.07373046875"